OpenStudy (anonymous):

The sum of three consecutive integers is 267. What is the largest integer. plz work it out so I can see how to do it

OpenStudy (midhun.madhu1987):

Let the 3 numbers be: x, x+1 and x+2 Then x+x+1+x+2 = 267 find x

OpenStudy (anonymous):

so whats the answer

OpenStudy (midhun.madhu1987):

3x+3=267 3x=267-3 3x=264 x=264/3 = 88 so the numbers are 88, 89 and 90

OpenStudy (anonymous):

thanks